# An algorithmic version of the Hajnal–Szemerédi theorem

### v2

@@ 6            Choose $T\in U$ such that $S^{+}(Q\cup\{T\})\leq t_{I}+2(r-1)$;
We choose the sub-family $K^{\prime}$ of $K$ by repeated applications of the pigeonhole principle. Indeed, we track the sum of the absolute values of the entries of the slack vectors restricted to $I$. More precisely, for any $K_{r}$-tiling $Q$ on $V$, let $\textbf{s}_{I}(Q):=\textbf{s}(Q)|_{I}$. Denote by $\textbf{s}_{I,j}(Q)$ the $j$-th coordinate of $\textbf{s}_{I}(Q)$ with $j\in[|I|]$. Let $S^{+}(Q):=\sum_{j\in[|I|]}|\textbf{s}_{I,j}(Q)|$. Note that for any copy $T^{\prime}$ of $K_{r}$ in $K$, we know $S^{+}(\{T^{\prime}\})\leq 2(r-1)$. We shall let Algorithm 1 read the members of $K$ in a smooth way in the sense that the partial $K_{r}$-tiling $Q$ considered by Algorithm 1 will never have large $S^{+}(Q)$. When $Q$ is large enough, by the pigeonhole principle we can find a nonempty $K_{r}$-tiling $Q^{\prime}$ with $\textbf{s}_{I}(Q^{\prime})=\textbf{0}$. We then remove the members of $Q^{\prime}$, put the members of $Q\setminus Q^{\prime}$ back to $U$ and repeat this process.
@@
The lemma is trivial if $(3t_{I})^{r}\geq n/r$. So we may assume that $(3t_{I})^{r}<n/r$. We execute Algorithm 1 and add copies of $K_{r}$ from $U$ to $Q$. If $S^{+}(Q)\leq t_{I}$, then $S^{+}(Q\cup\{T\})\leq t_{I}+2(r-1)$ for any $T\in U$, as $S^{+}(\{T\})\leq 2(r-1)$. First, since each time when we remove a $K_{r}$-tiling $Q^{\prime}$ (on lines 10-11), we have $\textbf{s}_{I}(Q^{\prime})=\textbf{0}$, at any stage of the algorithm, we have $\textbf{s}_{I}(Q\cup U)=\textbf{s}_{I}(K)$, and thus $S^{+}(Q\cup U)=S^{+}(K)=t_{I}$. Another immediate consequence is that for the final output $K^{\prime}$, as $U=\emptyset$ and $Q=K^{\prime}$, we conclude that $\textbf{s}_{I}(K^{\prime})=\textbf{s}_{I}(K)$.
@@
We next show that when $S^{+}(Q)>t_{I}$ the choice of $T$ as on line 4 is always possible. Indeed, suppose there is a moment that $S^{+}(Q)>t_{I}$ holds but it does not in the previous iteration. Then $t_{I}<S^{+}(Q)\leq t_{I}+2(r-1)$. Since $S^{+}(Q\cup U)=t_{I}<S^{+}(Q)$, there exists a copy $T$ of $K_{r}$ in $U$ such that $S^{+}(Q\cup\{T\})<S^{+}(Q)\leq t_{I}+2(r-1)$, that is, we can always choose $T\in U$ as on line 4 of Algorithm 1. Hence throughout the process, we can guarantee $S^{+}(Q)\leq t_{I}+2(r-1)$.
@@
It remains to show that when $Q$ gets large enough, on line 10 we can find the desired $Q^{\prime}$, which will also upper bound $|K^{\prime}|$. Note that each time we execute line 11, we “restart” the process by putting the copies of $K_{r}$ in $Q\setminus Q^{\prime}$ back to $U$. Therefore, let $Q$ be a nonempty ordered set of copies of $K_{r}$ before the execution of line 10, and for $i\in[|Q|]$ let $Q_{i}$ be the set of the first $i$ copies of $K_{r}$ in $Q$. Then for each $i\in[|Q|]$, by line 4, $Q_{i}$ satisfies that $S^{+}(Q_{i})\in[0,t_{I}+2(r-1)]$. Hence $\textbf{s}_{I,j}(Q_{i})\in[-(t_{I}+2(r-1)),t_{I}+2(r-1)]$ for each $i\in[|Q|]$ and $j\in[|I|]$. Therefore, the number of possible values of $\textbf{s}_{I}(Q_{i})$ is at most $(2(t_{I}+2(r-1))+1)^{|I|}\leq(3t_{I})^{r}$. Moreover, if line 11 is not executed in this iteration of the while loop, then in particular $\textbf{s}_{I}(Q_{i})\neq\textbf{s}_{I}(Q_{j})$ for any $1\leq i<j\leq|Q|$, we conclude that $|Q|\leq(3t_{I})^{r}$. In particular, we have $|K^{\prime}|\leq(3t_{I})^{r}$. ∎
@@
Next we give our color-coding lemma, which slightly extends its original formulation. Given a partition $(V_{1},V_{2},\dots,V_{r^{\prime}})$ of $V(G)$, consider all types of $K_{r}$, that is, all distinct vertex distributions of $K_{r}$ on $V(G)$. Recall that there are $y={{r+r^{\prime}-1}\choose{r^{\prime}-1}}$ different types of $K_{r}$. We fix an arbitrary ordering of these $y$ types. Given a $K_{r}$-tiling of size $m$, its pattern is an $m$-dimension vector where each digit is in $[y]$ representing its type. Denote by $\mathcal{P}_{m}$ the collection of all patterns of $K_{r}$-tilings of size $m$.
@@ 6. Extremal case
In this section we study the extremal case. Throughout the proof, we only consider
$c=o(n^{1/r^{r+3}}),$
as otherwise, we can check all possibilities by brute force, which takes time $O(n!)=2^{O(n\log n)}=2^{O(c^{r^{r+4}})}<n^{O(1)}2^{O(c^{r^{r+4}})}$.
@@ 6.1. Find sparse sets
Suppose $r\geq 3$ and $1/n\ll\gamma\ll\mu\ll\lambda\ll 1/r$. Let $G$ be an $n$-vertex graph with $\delta(G)\geq(1-1/r)n-c$. If $G$ has at least one $\gamma$-sparse $\frac{n}{r}$-set, then by Algorithm 2 we can find $s$ pairwise disjoint $\mu$-sparse $\frac{n}{r}$-sets $A_{1},A_{2},\ldots,A_{s}$ of $G$ in time $O(n^{2.376})$ such that $B:=V(G)\setminus\cup_{i=1}^{s}A_{i}$ has no $\lambda$-sparse $\frac{n}{r}$-set of $G$, where $1\leq s\leq r$.
@@
Let $s$ be the number of sparse sets that Algorithm 2 outputs. Let $\mu=\mu_{s}$ and $\lambda=\gamma_{s+1}$. Suppose that Algorithm 2 halts when $i\leq r-1$. Then $s=i-1$ and we can find $s$ pairwise disjoint $\mu$-sparse $\frac{n}{r}$-sets $A_{1},A_{2},\ldots,A_{s}$ such that $B:=G\setminus\cup_{j=1}^{s}A_{j}$ has no $\lambda$-sparse $\frac{n}{r}$-set.
Otherwise suppose that Algorithm 2 halts with $i=r$. Then either $s=r-1$ or $s=r$. If $s=r-1$, then we can find $r-1$ pairwise disjoint $\mu$-sparse $\frac{n}{r}$-sets $A_{1},A_{2},\ldots,A_{r-1}$ such that $B:=G\setminus\cup_{j=1}^{r-1}A_{j}$ is not a $\lambda$-sparse $\frac{n}{r}$-set. If $s=r$, then we can find $r$ pairwise disjoint $\mu$-sparse $\frac{n}{r}$-sets $A_{1},A_{2},\ldots,A_{r}$ of $G$.
@@
We need the following result when $s=r-2$ and $B$ is close to the union of two disjoint cliques.
@@
Recall that by Lemma 6.1, we have found $s$ pairwise disjoint $\mu$-sparse $\frac{n}{r}$-sets $A_{1},A_{2},\ldots,A_{s}$ of $G$ such that $B:=V(G)\setminus\cup_{i=1}^{s}A_{i}$ has no $\lambda$-sparse $\frac{n}{r}$-set of $G$, where $1/n\ll\mu\ll\lambda\ll 1/r$. In the rest of the proof, there is a tiny difference between the cases $11\leq s\leq r-$ 1 and $s=r$. We will first prove the former case with full details in Sections 6.3-6.6, and explain how to modify the approach to accommodate the latter in Section 6.7.

### v3

@@ 6     Choose $T\in U$ such that $S^{+}(Q\cup\{T\})\leq t_{I}+2(r-1)$;
In the proof of Lemma 1.8, we choose the sub-family $K^{\prime}$ of $K$ by repeated applications of the pigeonhole principle. Indeed, we track the sum of the absolute values of the entries of the slack vectors restricted to $I$. More precisely, for any $K_{r}$-tiling $Q$ on $V$, let $\textbf{s}_{I}(Q):=\textbf{s}(Q)|_{I}$. Denote by $\textbf{s}_{I,j}(Q)$ the $j$-th coordinate of $\textbf{s}_{I}(Q)$ with $j\in[|I|]$. Let $S^{+}(Q):=\sum_{j\in[|I|]}|\textbf{s}_{I,j}(Q)|$. Note that for any copy $T^{\prime}$ of $K_{r}$ in $K$, we know $S^{+}(\{T^{\prime}\})\leq 2(r-1)$. We shall let Algorithm 1 read the members of $K$ greedily but with a slack constraint in the sense that the partial $K_{r}$-tiling $Q$ considered by Algorithm 1 will never have large $S^{+}(Q)$. When $Q$ is large enough, by the pigeonhole principle we can find a nonempty $K_{r}$-tiling $Q^{\prime}$ with $\textbf{s}_{I}(Q^{\prime})=\textbf{0}$. We then remove the members of $Q^{\prime}$, put the members of $Q\setminus Q^{\prime}$ back to $U$ and repeat this process.
@@
We first prove the case $r\geq 4$. The lemma is trivial if $(3t_{I})^{r}\geq n/r$. So we may assume that $(3t_{I})^{r}<n/r$. We execute Algorithm 1 and add copies of $K_{r}$ from $U$ to $Q$. If $S^{+}(Q)\leq t_{I}$, then $S^{+}(Q\cup\{T\})\leq t_{I}+2(r-1)$ for any $T\in U$, as $S^{+}(\{T\})\leq 2(r-1)$. First, since each time when we remove a $K_{r}$-tiling $Q^{\prime}$ (on lines 10-11), we have $\textbf{s}_{I}(Q^{\prime})=\textbf{0}$, at any stage of the algorithm, we have $\textbf{s}_{I}(Q\cup U)=\textbf{s}_{I}(K)$, and thus $S^{+}(Q\cup U)=S^{+}(K)=t_{I}$. Another immediate consequence is that for the final output $K^{\prime}$, as $U=\emptyset$ and $Q=K^{\prime}$, we conclude that $\textbf{s}_{I}(K^{\prime})=\textbf{s}_{I}(K)$.
@@
We next show that when $S^{+}(Q)>t_{I}$ the choice of $T$ as on line 4 of Algorithm 1 is always possible. Indeed, suppose there is a moment that $S^{+}(Q)>t_{I}$ holds but it did not hold in the previous iteration. Then $t_{I}<S^{+}(Q)\leq t_{I}+2(r-1)$. Since $S^{+}(Q\cup U)=t_{I}<S^{+}(Q)$, there exists a copy $T$ of $K_{r}$ in $U$ such that $S^{+}(Q\cup\{T\})<S^{+}(Q)\leq t_{I}+2(r-1)$, that is, we can always choose $T\in U$ as described in line 4 of Algorithm 1. Hence throughout the process, we can guarantee $S^{+}(Q)\leq t_{I}+2(r-1)$.
@@
It remains to show that when $Q$ gets large enough, on line 10 we can find the desired $Q^{\prime}$, which will also upper bound $|K^{\prime}|$. Note that each time we execute line 11, we “restart” the process by putting the copies of $K_{r}$ in $Q\setminus Q^{\prime}$ back to $U$. Therefore, let $Q$ be a nonempty ordered set of copies of $K_{r}$ before the execution of line 10, and for $i\in[|Q|]$ let $Q_{i}$ be the set of the first $i$ copies of $K_{r}$ in $Q$. Then for each $i\in[|Q|]$, by line 4, $Q_{i}$ satisfies that $S^{+}(Q_{i})\in[0,t_{I}+2(r-1)]$. Hence
  • ($\dagger$)
    $\textbf{s}_{I,j}(Q_{i})\in[-(t_{I}+2(r-1)),t_{I}+2(r-1)]$ for each $i\in[|Q|]$ and $j\in[|I|]$.
Therefore, the number of possible values of $\textbf{s}_{I}(Q_{i})$ is at most $(2(t_{I}+2(r-1))+1)^{|I|}\leq(3t_{I})^{r}$. Moreover, if line 11 is not executed in this iteration of the while loop, then in particular $\textbf{s}_{I}(Q_{i})\neq\textbf{s}_{I}(Q_{j})$ for any $1\leq i<j\leq|Q|$, we conclude that $|Q|\leq(3t_{I})^{r}$. In particular, we have $|K^{\prime}|\leq(3t_{I})^{r}$.
Finally, let $r=3$ and we shall obtain a refined bound. First if $r^{\prime}=2$, then by ($\dagger$), the number of possible values of $\textbf{s}_{I}(Q_{i})$ is at most $3t_{I}$. Indeed, this is trivial if $|I|=1$; for $|I|=2$ it also holds as $\textbf{s}_{I,1}(Q_{i})=-\textbf{s}_{I,2}(Q_{i})$ and there are at most $3t_{I}$ choices for either of them.
Now suppose $r^{\prime}=3$. Again the result is trivial if $|I|=1$. If $|I|\geq 2$, then note that $t_{[3]}\leq 2t_{I}$, and thus it suffices to show that the $K_{r}$-tiling $K^{\prime}$ can be chosen to have size at most $4t_{[3]}$. Indeed, if $|I|=2$, then by the assumption, we obtain a $K_{r}$-tiling $K^{\prime}$ of size at most $4t_{[3]}$ such that $\textbf{s}(K^{\prime})=\textbf{s}(K)$. Hence $|K^{\prime}|\leq 4t_{[3]}\leq 8t_{I}$ and $\textbf{s}(K^{\prime})|_{I}=\textbf{s}(K)|_{I}$, that is, $K^{\prime}$ is a desired $K_{r}$-tiling.
For $I=[3]$, note that there are three types of index vectors of triangles up to permutation, that is, $(3,0,0)$, $(2,1,0)$ and $(1,1,1)$, whose slack vectors are (2,1,1)(2,-1,-1), (1,0,1)(1,0,-1) and $(0,0,0)$, respectively. Suppose $K^{\prime}$ is minimal subject to the condition $\textbf{s}(K^{\prime})=\textbf{s}(K)$. Therefore, $K^{\prime}$ contains no triangles of index vector $(1,1,1)$, as we may freely remove them without changing $\textbf{s}(K^{\prime})$. We now notice a useful fact stated as follows.
Claim 5.3.
If there is $i\in[3]$ such that all slack vectors of triangles in $K^{\prime}$ have non-negative (non-positive) values on the $i$-th digit, then $|K^{\prime}|\leq 4t_{I}$.
Proof.
Suppose all triangles of $K^{\prime}$ have slack vectors non-positive in the first digit. Let $K_{1}^{\prime}\subseteq K^{\prime}$ be the triangles in $K^{\prime}$ whose first digit negative. Then we have $|K_{1}^{\prime}|\leq t_{I}$. For $K_{2}^{\prime}:=K^{\prime}\setminus K_{1}^{\prime}$, note that all triangles in it have the same slack vectors, either all (0,1,1)(0,1,-1), or all (0,1,1)(0,-1,1). Without loss of generality, suppose all of them have slack (0,1,1)(0,1,-1), yielding that $\textbf{s}(K_{2}^{\prime})=(0,|K_{2}^{\prime}|,-|K_{2}^{\prime}|)$. As $|\textbf{s}_{I,j}(K_{1}^{\prime})|\leq 2|K_{1}^{\prime}|\leq 2t_{I}$ for $j=2,3$, we conclude that $||K_{2}^{\prime}|-2t_{I}|\leq t_{I}$, implying $|K_{2}^{\prime}|\leq 3t_{I}$. Therefore, we have $|K^{\prime}|\leq 4t_{I}$. ∎
Now we split the discussion according to the types of triangles in $K^{\prime}$. Let $\mathcal{I}$ be the set of the slacks of triangles in $K^{\prime}$, that is, $\mathbf{i}\in\mathcal{I}$ if and only if $K^{\prime}$ contains a triangle of slack $\mathbf{i}$.
Case 1. $K^{\prime}$ contains triangles of slack vectors (2,1,1)(2,-1,-1) and (1,2,1)(-1,2,-1), that is, (2,1,1)(2,-1,-1), $(-1,2,-1)\in\mathcal{I}$. We may suppose that (1,0,1)(-1,0,1) or (0,1,1)(0,-1,1) is in $\mathcal{I}$, as otherwise we conclude $|K^{\prime}|\leq 4t_{I}$ by Claim 5.3. Without loss of generality, suppose $(-1,0,1)\in\mathcal{I}$. Then by minimality, we know that $K^{\prime}$ has no triangle with slack (1,0,1)(1,0,-1), (1,1,0)(-1,1,0) or (0,1,1)(0,-1,1) and at most one triangle with slack (1,1,0)(1,-1,0). Therefore, $K^{\prime}$ contains triangles of slack (2,1,1)(2,-1,-1), (1,2,1)(-1,2,-1), (1,0,1)(-1,0,1) and possibly (1,1,0)(1,-1,0) (at most one) and (0,1,1)(0,1,-1). Note that only (1,0,1)(-1,0,1) has positive slack on its third coordinate. Now if $K^{\prime}$ contains at most $t_{I}+3$ triangles of slack (1,0,1)(-1,0,1), then it contains at most $2t_{I}+3$ triangles of slack being one of (2,1,1)(2,-1,-1), (1,2,1)(-1,2,-1) and (0,1,1)(0,1,-1), yielding $|K^{\prime}|\leq 3t_{I}+7\leq 4t_{I}$. Otherwise $K^{\prime}$ contains at least $t_{I}+4$ triangles of slack (1,0,1)(-1,0,1), and thus contribute $t_{I}-4$ to the first digit of $\textbf{s}(K^{\prime})$. As $K^{\prime}$ contains at most one triangle of slack (1,1,0)(1,-1,0), this implies that $K^{\prime}$ contains at least two triangles of slack (2,1,1)(2,-1,-1). However, noticing the equation 2(2,1,1)+(1,2,1)+3(1,0,1)=(0,0,0)2(2,-1,-1)+(-1,2,-1)+3(-1,0,1)=(0,0,0), we obtain a contradiction with the minimality of $K^{\prime}$.
Case 2. $(2,-1,-1)\in\mathcal{I}$ and $(-1,2,-1),(-1,-1,2)\notin\mathcal{I}$. Similar to the previous case, by Claim 5.3, we may suppose that $(-1,1,0)\in\mathcal{I}$ ($\mathcal{I}$ needs to have (1,1,0)(-1,1,0) or (1,0,1)(-1,0,1), which are the only slack vectors with negative first coordinate). Then by minimality of $K^{\prime}$, we have $(1,-1,0),(-1,0,1)\notin\mathcal{I}$. Now, considering the third coordinate, we must have $(0,-1,1)\in\mathcal{I}$ by Claim 5.3 and so $(0,1,-1),(1,0,-1)\notin\mathcal{I}$ by minimality. In summary, $\mathcal{I}=\{(2,-1,-1),(-1,1,0),(0,-1,1)\}$ and $K^{\prime}$ contains at most one triangle of slack (1,1,0)(-1,1,0). Note that (1,1,0)(-1,1,0) is the only one with positive coordinate on the second coordinate and $K^{\prime}$ has at most one such triangle. Then as the other two have negative second coordinate, we conclude that $|K^{\prime}|\leq t_{I}+2$.
Case 3. $(2,-1,-1),(-1,2,-1),(-1,-1,2)\notin\mathcal{I}$. Clearly in this case $|\mathcal{I}|\leq 3$. If $|\mathcal{I}|\leq 2$, then by Claim 5.3, $|K^{\prime}|\leq 4t_{I}$. Otherwise $|\mathcal{I}|=3$, and without loss of generality, suppose $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}\in\mathcal{I}$ (they are all permutations of (1,1,0)(-1,1,0)). We indeed have $\mathbf{v}_{3}=\mathbf{v}_{1}+\mathbf{v}_{2}$ and for any non-zero coordinate of $\mathbf{v}_{3}$, the equation $\mathbf{v}_{3}=\mathbf{v}_{1}+\mathbf{v}_{2}$ restricts on this coordinate is either $1=0+1$ or 1=0+(1)-1=0+(-1). Therefore, we conclude $|K^{\prime}|\leq 4t_{I}$ by Claim 5.3.
We conclude the proof by symmetry. ∎
@@
Next we give our color-coding lemma, which slightly extends its original formulation. Given a partition $(V_{1},V_{2},\dots,V_{r^{\prime}})$ of $V(G)$, consider all types of $K_{r}$, that is, all distinct vertex distributions of $K_{r}$ on $V(G)$. Recall that there are $y={{r+r^{\prime}-1}\choose{r^{\prime}-1}}$ different types of $K_{r}$. We fix an arbitrary ordering of these $y$ types. Given a $K_{r}$-tiling of size $m$, its pattern is an $m$-dimensional vector where each digit is in $[y]$ representing its type. Denote by $\mathcal{P}_{m}$ the collection of all patterns of $K_{r}$-tilings of size $m$.
@@ 6. Extremal case
In this section we study the extremal case. Throughout the proof, it suffices to deal with
$c=\begin{cases}o(n^{1/r^{r+3}}),\text{ if }r\geq 4,\\ o(n),\hfill\text{ if }r=3,\end{cases}$
as otherwise, Theorem 1.4 implies that we can determine whether $G$ contains a $K_{r}$-factor in time $O(n^{r}+2^{O(n)})=2^{O(n)}$, which is $2^{O(c)}$ for $r=3$ and $2^{O(c^{r^{r+3}})}$ for $r\geq 4$, as desired. Moreover, throughout this section, we work with the following parameters satisfying the hierarchy
$1/n\ll\gamma\ll\mu\ll\lambda\ll 1/r,$
which we will omit in the subsequent statements of lemmas.
@@ 6.1. Find sparse sets
Let $G$ be an $n$-vertex graph with $\delta(G)\geq(1-1/r)n-c$. If $G$ has at least one $\gamma$-sparse $\frac{n}{r}$-set, then by Algorithm 2 we can find $s$ pairwise disjoint $\mu$-sparse $\frac{n}{r}$-sets $A_{1},A_{2},\ldots,A_{s}$ of $G$ in time $O(n^{2.376})$ such that $B:=V(G)\setminus\cup_{i=1}^{s}A_{i}$ has no $\lambda$-sparse $\frac{n}{r}$-set of $G$, where $1\leq s\leq r$.
@@
Let $s$ be the number of sparse sets that Algorithm 2 outputs. Let $\mu=\mu_{s}$ and $\lambda=\gamma_{s+1}$. Suppose that Algorithm 2 halts when $i\leq r-1$. Then s=i1s=i-1 and we can find $s$ pairwise disjoint $\mu$-sparse $\frac{n}{r}$-sets $A_{1},A_{2},\ldots,A_{s}$ such that $B:=G\setminus\cup_{j=1}^{s}A_{j}$ has no $\lambda$-sparse $\frac{n}{r}$-set.
Otherwise suppose that Algorithm 2 halts with $i=r$. Then either s=r1s=r-1 or $s=r$. If s=r1s=r-1, then we can find r1r-1 pairwise disjoint $\mu$-sparse $\frac{n}{r}$-sets $A_{1},A_{2},\ldots,A_{r-1}$ such that $B:=G\setminus\cup_{j=1}^{r-1}A_{j}$ is not a $\lambda$-sparse $\frac{n}{r}$-set. If $s=r$, then we can find $r$ pairwise disjoint $\mu$-sparse $\frac{n}{r}$-sets $A_{1},A_{2},\ldots,A_{r}$ of $G$.
@@
We need the following result when s=r2s=r-2 and $B$ is close to the union of two disjoint cliques.
@@
Recall that by Lemma 6.1, we have found $s$ pairwise disjoint $\mu$-sparse $\frac{n}{r}$-sets $A_{1},A_{2},\ldots,A_{s}$ of $G$ such that $B:=V(G)\setminus\cup_{i=1}^{s}A_{i}$ has no $\lambda$-sparse $\frac{n}{r}$-set of $G$, where $1/n\ll\mu\ll\lambda\ll 1/r$. In the rest of the proof, there is a tiny difference between the cases $1\leq s\leq r-1$ and $s=r$. We will first prove the former case with full details in Sections 6.2-6.6, and explain how to modify the approach to accommodate the latter in Section 6.7.